OpenStudy (anonymous):

sole this system please put answer as ordered pair x+y=-3 2x+3y=-11

OpenStudy (anonymous):

First, isolate x or y in the first equation.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

Because they're both positive, you just subtract one. For x, the equation would end up x=3-y

OpenStudy (anonymous):

Follow?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok. Now that x is isolated, we can replace each x in the other equation with 3-y, because we now know that is what x equals. That gives us an equation of 2(3-y)+3y=-11

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

yeah im with you

OpenStudy (anonymous):

Good! Now we need to isolate y in our new equation. Do you know how to do that?

OpenStudy (anonymous):

y=2x+-11

OpenStudy (anonymous):

any help?

OpenStudy (anonymous):

Sorry for the delay!! I was helping somebody else.

OpenStudy (anonymous):

Oops, you solved for y without 3-y in the new equation.

OpenStudy (anonymous):

Because there's only one variable in the equation, you should be able to come up with a value for y without another variable in it.

OpenStudy (anonymous):

not understanding

OpenStudy (anonymous):

The equation that we came up with for the second equation was 2(3-y)+3y=-11

OpenStudy (anonymous):

So now we need to solve for y in that equation

OpenStudy (anonymous):

6-2y+3y=-11 6-5y=-11 +5 +5 6=-6 /6 /6 y=-1?

OpenStudy (anonymous):

-2y+3y=y, not -5y

OpenStudy (anonymous):

But good first effort. I'm honestly ecstatic that you're trying to do the work yourself, rather than relying purely on me :)

OpenStudy (anonymous):

dang that was my bad

OpenStudy (anonymous):

Haha it's fine

OpenStudy (anonymous):

im really trying to learn

OpenStudy (anonymous):

I can see that! I'm happy to help teach you as best I can.

OpenStudy (anonymous):

6-2y+3y=-11 6-5y=-11 +5 +5 6=-6 /6 /6 y=-1?

OpenStudy (anonymous):

Isn't that the same thing?

OpenStudy (anonymous):

Or my internet connection is messed up. That's also a possibility

OpenStudy (anonymous):

no i didnt type the same thing

OpenStudy (anonymous):

Hmmmmm- would you mind copy/pasting your latest work then and repost it?

OpenStudy (anonymous):

computer wont let me

OpenStudy (anonymous):

what was your final answer?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Well, either way, that's still incorrect. Let me show you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2(3-y)+3y=-11 Distribute 6-2y+3y=-11 Add Like Terms 6+y=-11 Subtract 6 from both sides. y=-17

OpenStudy (anonymous):

Now then- That's not your final answer, but we're very close. Now that we know what y is, we replace each y in the first equation with -17. This gives you an equation of x+(-17)=-3

OpenStudy (anonymous):

This can be rewritten as x-17=-3 Now you just simplify by adding 17 to both sides of the equation, for a solution of x= 14

OpenStudy (anonymous):

So your final answer is x=14 and y=-17

OpenStudy (anonymous):

Well, I have to go now, but good luck with your math!!